6. The manufacturer of high-quality flatbed scanners is trying to decide what price to set for its product. The costs of production and the demand for the product are assumed to be as follows:

TC = 500,000 + 0.85Q + (0.015Q)(0.015Q)

Q = 14,166 - 16.6P

a. Determine the short-run profit-maximizing price.

b. Plot this information on a graph showing AC, AVC, MC, P, and MR.
Answer #1

TC = 500,000 + 0.85Q + (0.015Q)2 = 500,000 + 0.85Q + 0.000225Q2

MC = dTC/dQ = 0.85 + 0.00045Q

Q = 14,166 - 16.6P, so

16.6P = 14,166 - Q

P = (14,166 - Q) / 16.6

(a)

Profit is maximized when MR = MC

Total revenue (TR) = P x Q = (14,166Q - Q2) / 16.6

MR = dTR/dQ = (14,166 - 2Q) / 16.6

(14,166 - 2Q) / 16.6 = 0.85 + 0.00045Q

14,166 - 2Q = 14.11 + 0.00747Q

2.00747Q = 14,151.89

Q = 7,050

P = (14,166 - 7,050) / 16.6 = 7,116 / 16.6 = 428.67

(b)

AC = TC/Q = (500,000/Q) + 0.85 + 0.000225Q

TVC = 0.85Q + 0.000225Q2

AVC = TVC/Q = 0.85 + 0.000225Q

Data table used:

Q P MR AC AVC MC
0 853.37 853.37
1000 793.13 732.89 501.08 1.075 1.3
2000 732.89 612.41 251.30 1.3 1.75
3000 672.65 491.93 168.19 1.525 2.2
4000 612.41 371.45 126.75 1.75 2.65
5000 552.17 250.96 101.98 1.975 3.1
6000 491.93 130.48 85.53 2.2 3.55
7000 431.69 10.00 73.85 2.425 4
8000 371.45 -110.48 65.15 2.65 4.45
9000 311.20 -230.96 58.43 2.875 4.9
10000 250.96 -351.45 53.10 3.1 5.35

Graph: